# What Happens During the Loss Plateau? Understanding Abrupt Learning in Transformers

### v1

@@ 1 Introduction
Training Transformers on mathematical or algorithmic tasks often exhibits an intriguing “abrupt learning” phenomenon in their training dynamics, where the model’s performance plateaus at a suboptimal level for an extended period before suddenly and rapidly converging to the optimal solution nanda2023progress ; barak2022hidden ; singh2024 ; zhang2025incontext ; wang2025rich (Figures 1 and 2). This is often considered an example of the broader phenomenon of “emergence,” where model capabilities appear to arise discontinuously and unpredictably with increasing amount of parameters, training data, or training steps wei2022emergent . Understanding these sharp phase transitions in learning trajectories is crucial for gaining deeper insights into how Transformer models learn and develop their sophisticated capabilities.
Despite recent progress in understanding such abrupt learning dynamics in Transformers for specific tasks like in-context learning singh2024 ; zhang2025incontext ; wang2025rich ; chen2024trainingdynamics ; reddy2023mechanisticbasisdatadependence , parity learning barak2022hidden , Markov chains edelman_evolution , grammar learning percolation , and matrix completion gopalani2024abrupt , a unifying account of the model evolution during loss plateau is still missing. Further, many of these works require specific assumptions on model and data that limit their generality.
@@
We identify novel implicit biases that underlie the early plateau period of Transformer training: the model learns a partial solution while being biased toward degenerate patterns in its outputs and internal representations. We further study the pivotal role of attention map learning in driving these phenomena and overcoming the performance plateau. See Figure 1 for an overview of our findings. Our specific contributions are:
  • Partial solutions during plateau: We show that during the initial loss plateau, the model often learns a partial solution, which correctly predicts a subset of easier tokens within a sequence—those that might be intuitively simpler to learn (e.g., copying the first element in a moving-window sum, or predicting the final carry-over in multi-digit addition)—while failing on more complex parts of the task. This pattern is observed across diverse algorithmic tasks (Table 1).
  • Repetition bias in outputs: We identify a strong repetition bias during the plateau, where the model tends to output repetitive tokens. This bias can be quantified by metrics such as a direct count of repeated subsequent tokens or the entropy of the output token distribution. Such repetitions significantly increase during the early training steps and then markedly decrease as the performance starts to improve (Figure 2(b)).
  • Internal representation collapse: The output repetition bias is accompanied by representation collapse, where hidden representations for different tokens become nearly parallel (e.g., cosine similarity often exceeds $0.9$), indicating a degenerate representational geometry inside Transformers. Subsequently, this representational similarity drops significantly as the model’s performance improves, signifying a diversification of internal representations (Figure 2(b)).
  • Crucial role of attention map learning: We find that gradual learning of the optimal attention pattern can commence during the loss plateau, before the sudden drop in loss (Figure 2(b)). By directly intervening on the attention map during training—for instance, by biasing the attention scores towards or away from the optimal configuration—we can observe tangible changes in the duration of loss plateau and the severity of degenerate behaviors like repetition bias and representation collapse.
@@ 2 Setup, Abrupt Learning, and Attention Map
We mainly present results for the moving-window-sum task in the main text, which we define below; we also validate our findings on various other algorithmic tasks like multi-digit addition, permutations, histogram, prefix-sum, etc., in Appendix B.
@@
Here, $x_{1},\ldots,x_{n}$ are the input sequence, ${\rm SEP}$ is a separator token, and the task is to complete the sequence with outputs $y_{1},\ldots,y_{n}$. In the experiments in the main paper, we use $n=16$, $p=17$, $x_{i}\sim{\rm Unif}\{1,2,\ldots,16\}$ and ${\rm SEP}=17$. We denote the full vocabulary $V:=\{0,1,2,\ldots,17\}.$
@@
${\rm TF}_{\theta}(s_{1},s_{2},\ldots,s_{L})={\rm LM}\circ{\rm(Id+MLP)}\circ{\rm(Id+Attn)}\circ{\rm Embed}(s_{1},s_{2},\ldots,s_{L})$
@@
$[{\rm Attn}(h_{1},h_{2},\ldots,h_{L})]_{i}=W_{O}\left(\sum_{j=1}^{i}(h_{j}^{\top}W_{K}^{\top}W_{Q}h_{i})W_{V}h_{j}\right);\quad W_{O},W_{K},W_{Q},W_{V}\in\mathbb{R}^{d\times d}.$
${\rm MLP}$ denotes the 2-layer neural net $h_{i}\mapsto W_{2}(\sigma(W_{1}h_{i}))$ for $W_{2}\in\mathbb{R}^{d\times 4d},W_{1}\in\mathbb{R}^{4d\times d},$ and $\sigma$ the GELU activation. ${\rm LM}$ is a linear layer that maps the hidden state $h_{i}\in\mathbb{R}^{d}$ to logits $v_{i}\in\mathbb{R}^{\left|V\right|}$. Note that all linear maps above implicitly include a bias term, and we use pre-LayerNorm so that before the Attn, MLP, and LM, a LayerNorm operation is applied to the hidden states $h_{i}$. For generating sequences, we use greedy decoding i.e. output token is determined by the maximum logit over the vocabulary (please see Appendix A for more implementation details). We use linear attention to avoid vanishing gradient issues from softmax attention being a contributing factor toward abrupt learning, which was argued in eureka . We also show similar results on softmax attention, multi-layer / multi-head models, and models with varying $d$ in Appendix D.
@@
The model is trained to minimize the standard next-token-prediction cross-entropy loss over the full sequence i.e. $(x_{1},\ldots,x_{n},{\rm SEP},y_{1},\ldots y_{n})$ for the MWS task. We evaluate accuracy over the output portion of the sequence, i.e., $y_{1},\ldots,y_{n}$, averaged over these $n$ positions. We use the Adam optimizer with a constant learning rate $10^{-4}$ and no weight decay. The training is conducted in an online / single-epoch fashion, where a new batch of 256 training samples is drawn from the data distribution at each training step. Note that in this setup, the training and test losses essentially coincide. For completeness, we also show similar results on the SGD optimizer in Appendix D.
@@
Following the training procedure described above will result in a characteristic abrupt learning curve, where the training/test loss is stuck at some sub-optimal value for a significant number of steps, before suddenly and rapidly decreasing to its optimal value (Figure 2(a)). This drop in loss is accompanied by a similarly rapid increase in accuracy, indicating that the optimal solution is learned abruptly.
@@
We analyze the attention map at different points during training. We find that the attention map shows a sparse, interpretable pattern after the sudden loss drop, while no such pattern is shown before the sudden drop (Figure 1). For the MWS task, this optimal attention pattern corresponds to each output token $y_{i}$ attending only to the input tokens relevant to its computation, i.e., attending to $x_{1}$ for $y_{1},$ and to $x_{i},x_{i-1}$ for $y_{i},i\geq 2.$ We further use an Attention Progress Measure (APM) to record the progress of the attention map toward its optimal pattern during training, defined as
${\rm APM}:=\frac{\sum_{(i,j)\in\Omega}|A_{ij}|}{\sum_{(i,j)}|A_{ij}|},$
where $A_{ij}$ denotes the attention score allocated to the $j^{th}$ token when computing output at the $i^{th}$ position in the sequence, and $\Omega$ is the set of position pairs in the optimal attention map. This measure is defined with absolute values due to our choice of linear attention so that $A_{ij}$ could be positive or negative. In experiments, we calculate ${\rm APM}$ averaged over a random batch of sequences.
Figure 2(b) shows that the ${\rm APM}$ monotonically increases from near 00 to near $0.8$ during training, and its increase is more gradual than the loss/accuracy dynamics. In particular, ${\rm APM}$ already increases to a nontrivial value during the loss plateau and before the sudden loss drop.
@@ 3 Implicit Biases in the Early Phase of Training
During the loss plateau, the model often has already learned to implement a partial solution to the task. This means it correctly predicts a subset of the output tokens, typically those corresponding to an intuitively simpler part of the problem, while failing on the more complex parts. For instance, in the MWS task, the model quickly learns to predict the first output token $y_{1}$ correctly (see Figure 2(a) for the first-token accuracy), as it is simply a copy of the first input token $x_{1}$, while the overall loss remains high and accuracy on subsequent tokens is poor. This ability to solve easier sub-components of the task early on is observed across various algorithmic problems (see Table 1 in Appendix B).
@@
$\rho:=\frac{1}{n-1}\sum_{i=1}^{n-1}\mathbf{1}[y_{i}=y_{i+1}].$
We observe that $\rho$ increases rapidly during the early phase of training, when the optimal attention map has not been learned yet (Figure 2(b)). Note that this frequency is small at initialization, and grows rapidly in the first $\approx 50$ steps to $\approx 0.8$, indicating that it is an implicit bias coming from gradient-based training.
@@
${\rm COS}_{i,j}:=\frac{\langle\mathbf{h}_{i},\mathbf{h}_{j}\rangle}{\|\mathbf{h}_{i}\|\|\mathbf{h}_{j}\|}$
where $\mathbf{h}_{i}\in\mathbb{R}^{d}$ is the hidden state at position $i$ in the sequence (this quantity is averaged over a random batch of sequences). We find that in the early phase of training, there is a rapid increase in ${\rm COS}_{i,j}$—averaged over all output positions $i,j$ except the first position, this quantity increases to $\approx 0.95$ (Figure 2(b)). Similar to repetitions, representation collapse is not present at initialization and only appears after a few steps of training. This is in contrast to the rank collapse phenomenon for deep softmax-attention Transformers anagnostidis2022signal that occurs at initialization. Also, while we focus on the final-layer representation (before the LM layer) in the main text, we show in Figure 28 that representation collapse happens in all intermediate layers to varying degrees.
@@ 4 The Role of Learning Attention
Observe that though the loss dynamics are abrupt, attention progress measure as well as repetitions and representation collapse are not (Figure 2(b)); that is, even when the loss is barely decreasing (between steps 50 and 150), attention progress measure notably increases, accompanied by a decrease in repetition frequency and representation collapse. Via training-time interventions, this section shows that learning the attention map plays a crucial role in shaping the loss plateau as well as repetitions and representation collapse.
@@
We start by verifying whether the attention layer is responsible for representation collapse during the early phase of training. To this end, we plot the cosine similarity of the residual stream for output tokens just before and after the attention layer. Formally, let the residual stream before attention layer (i.e., token + positional embeddings) be $\mathbf{h}_{i}\in\mathbb{R}^{d}$, and the residual stream after attention layer be $\mathbf{h}^{\prime}_{i}\in\mathbb{R}^{d}$, we measure the norm and pairwise cosine similarity for $\mathbf{h}_{i}$ and $\mathbf{h}^{\prime}_{i}$ in Figure 27.
@@
To study the role of attention map, we slightly modify the training process starting at different time points in training, biasing it towards (or away from) the optimal attention map to check if repetitions, representation collapse, and loss plateau are reduced (resp. amplified).
@@
We do the following: at training time, starting at step $t_{0},$ we multiply the attention map values for output tokens except the first position at $\Omega$ (i.e. optimal attention map positions) by a constant $c>0;$ for $c>1,$ this implies biasing the model towards the final (optimal) attention map, whereas for $0<c<1,$ this implies biasing the model away from the optimal attention map.
We find that, for $c>1$ and various values of $t_{0},$ such a scaling leads to lower average cosine similarity between hidden states, lower frequency of repetitions, and faster convergence (Figures 3 and 29). Whereas, for $0<c<1,$ we find the opposite: the model is in representation collapse state for a longer time and converges later compared to the non-scaled $(c=1)$ case, while the repetition frequency remains large throughout the plateau (Figure 4).
For example, for $t_{0}=0,c=10$, i.e. scaling 10×10$\times$10 × from the start of training, we find that the peak cosine similarity attained during training is $\approx 0.6$, much smaller than the $\approx 0.95$ attained for $c=1,$ and further the peak for $c=10$ is for negligible duration compared to that for $c=1$. Later values of $t_{0}=25,50,75$ show similar results wherein the cosine similarity drops immediately on the above biasing operation, followed by lower repetition frequency and convergence to optimal solution (Figure 29).
On the other hand, for $t_{0}=0,c=0.2,0.5$, the model takes much longer to converge and is in representation collapse / large repetition frequency state for much longer. This is in line with our expectation that lower attention map values for the optimal positions lead to slower learning and prolonged representation collapse.
@@
Training with Optimal Attention.
In this test, we initialize with the optimal attention map by fixing embeddings, LayerNorm for attention layer and attention layer weights to their final values at the end of a normal training run, so that at initialization, the correct attention map is already available to subsequent layers. We re-train the subsequent non-fixed layers starting from random initialization.
For the attention layer, we choose the set of parameters to initalize in 2 ways: (a) only Key, Query $(W_{K},W_{Q})$ weights, and (b) All of Key, Query, Value, Output $(W_{K},W_{Q},W_{V},W_{O})$ weights. We find that in both of these cases, learning only the subsequent layers (i.e. MLP, LM Head) take significantly shorter time than training the full model, without any significant representation collapse, repetitions or plateau in loss (Figure 5). Further, between (a) and (b), we find that additionally having $W_{O},W_{V}$ layers initialized to optimal values slightly speeds up learning, and average cosine similarity goes up to approx $0.15$ instead of $\approx 0.45$ when only initializing $W_{K},W_{Q}$ weights. This indicates that $W_{O},W_{V}$ layers also play a non-trivial role in causing representation collapse. This result confirms that attention map is a major bottleneck that leads to early representation collapse and loss plateau.
@@
Optimal MLP or Embeddings Do Not Help.
On the other hand, fixing MLP or embeddings (together with LM head) to their final optimal values and re-training the other components does not qualitatively change the training dynamics from the full training case, i.e., a significant loss plateau, repetition bias, and representation collapse still occur (Figure 5). This indicates that there is little benefit from having the optimal MLP or embeddings at initialization compared to attention map.

### v2

@@ 1 Introduction
Training Transformers on mathematical or algorithmic tasks often exhibits an intriguing “abrupt learning” phenomenon in their training dynamics, where the model’s performance plateaus at a suboptimal level for an extended period before suddenly and rapidly converging to the optimal solution [36, 3, 44, 54, 47] (Figures˜1 and 2). This is often considered an example of the broader phenomenon of “emergence,” where model capabilities appear to arise discontinuously and unpredictably with increasing amount of parameters, training data, or training steps [49]. Understanding these sharp phase transitions in learning trajectories is crucial for gaining deeper insights into how Transformer models learn and develop their sophisticated capabilities.
Despite recent progress in understanding such abrupt learning dynamics in Transformers for specific tasks like in-context learning [44, 54, 47, 10, 41, 15], parity learning [3], Markov chains [15], grammar learning [32, 9], and matrix completion [18], a unifying account of the model evolution during loss plateau is still missing. Further, many of these works require specific assumptions on model and data that limit their generality.
@@
We identify implicit biases that underlie the early plateau period of Transformer training: the model learns a partial solution while being biased toward degenerate patterns in its outputs and internal representations. We further study the pivotal role of attention map learning in driving these phenomena and overcoming the performance plateau. See Figure˜1 for an overview of our findings. Our specific contributions are:
  • Partial solutions during plateau: We show that during the initial loss plateau, the model often learns a partial solution, which correctly predicts a subset of easier tokens within a sequence—those that might be intuitively simpler to learn (e.g., copying the first element in a moving-window sum, or predicting the final carry-over in multi-digit addition)—while failing on more complex parts of the task. This pattern is observed across diverse algorithmic tasks (Table˜1).
  • Repetition bias in outputs: We identify a strong repetition bias during the plateau, where the model tends to output repetitive tokens. This bias can be quantified by metrics such as a direct count of repeated subsequent tokens or the entropy of the output token distribution. Such repetitions significantly increase during the early training steps and then markedly decrease as the performance starts to improve (Figure˜2(b)).
  • Internal representation collapse: The output repetition bias is accompanied by representation collapse, where hidden representations for different tokens become nearly parallel (e.g., cosine similarity often exceeds $0.9$), indicating a degenerate representational geometry inside Transformers. Subsequently, this representational similarity drops significantly as the model’s performance improves, signifying a diversification of internal representations (Figure˜2(b)).
  • Crucial role of attention map learning: We find that gradual learning of the optimal attention pattern can commence during the loss plateau, before the sudden drop in loss (Figure˜2(b)). By directly intervening on the attention map during training—for instance, by biasing the attention scores towards or away from the optimal configuration—we can observe tangible changes in the duration of loss plateau and the severity of degenerate behaviors like repetition bias and representation collapse.
@@ 2 Setup, Abrupt Learning, and Attention Map
We mainly present results for the moving-window-sum task in the main text, which we define below; we also validate our findings on various other algorithmic tasks like multi-digit addition, permutations, histogram, prefix-sum, etc., in Appendix˜B.
@@
Here, $x_{1},\ldots,x_{n}$ are the input sequence, $\rm SEP$} is a separator token, and the task is to complete the sequence with outputs $y_{1},\ldots,y_{n}$. In the experiments in the main paper, we use $n=16$, $p=17$, $x_{i}\sim{\rm Unif}\{1,2,\ldots,16\}$ and ${\rm SEP}=17$. We denote the full vocabulary $V:=\{0,1,2,\ldots,17\}$, and directly use these integers as token IDs when generating token embeddings for input to the model.
@@
${\rm TF}_{\theta}(s_{1},s_{2},\ldots,s_{L})={\rm LM}\circ{\rm(Id+MLP)}\circ{\rm(Id+Attn)}\circ{\rm Embed}(s_{1},s_{2},\ldots,s_{L})$ (1)
@@
${\rm Attn}(h_{1},h_{2},\ldots,h_{L})]_{i}=W_{O}\left(\sum_{j=1}^{i}(h_{j}^{\top}W_{K}^{\top}W_{Q}h_{i})W_{V}h_{j}\right);W_{O},W_{K},W_{Q},W_{V}\in\mathbb{R}^{d\times d}$ (2)
$\rm MLP$} denotes the 2-layer neural net $h_{i}\mapsto W_{2}(\sigma(W_{1}h_{i}))$ for $W_{2}\in\mathbb{R}^{d\times 4d},W_{1}\in\mathbb{R}^{4d\times d},$ and $\sigma$ the GELU activation. $\rm LM$} is a linear layer that maps the hidden state $h_{i}\in\mathbb{R}^{d}$ to logits $v_{i}\in\mathbb{R}^{\left|V\right|}$. Note that all linear maps above implicitly include a bias term, and we use pre-LayerNorm so that before the Attn, MLP, and LM, a LayerNorm operation is applied to the hidden states $h_{i}$. For generating sequences, we use greedy decoding i.e. output token is determined by the maximum logit over the vocabulary (please see Appendix˜A for more implementation details). We use linear attention to avoid small gradient issues from softmax function being a contributing factor toward abrupt learning, as argued in [20]. We also show similar results on softmax attention, multi-layer / multi-head models, and models with varying $d$ in Appendix˜F.
@@
The model is trained to minimize the standard next-token-prediction cross-entropy loss over the full sequence i.e. $(x_{1},\ldots,x_{n},{\rm SEP},y_{1},\ldots y_{n})$ for the MWS task. We evaluate accuracy over the output portion of the sequence, i.e., $y_{1},\ldots,y_{n}$, averaged over these $n$ positions. We use the Adam optimizer with a constant learning rate $10^{-4}$ and no weight decay. The training is conducted in an online / single-epoch fashion, where a new batch of 256 training samples is drawn from the data distribution at each training step. Note that in this setup, the training and test losses essentially coincide. For completeness, we verify in Appendix˜G that our reported phenomena occur when using different optimization algorithms (SGD, Muon [25]) or optimization hyperparameters (learning rate, LR schedules, batch size, initialization scale, weight decay) as well, and are not merely an artifact of a specific training setup. Code for experiments is available at github.com/pulkitgopalani/tf-loss-plateau.
@@
Following the training procedure described above will result in a characteristic abrupt learning curve, where the training/test loss is stuck at some sub-optimal value for a significant number of steps, before suddenly and rapidly decreasing to its optimal value (Figure˜2(a)). This drop in loss is accompanied by a similarly rapid increase in accuracy, indicating that the optimal solution is learned abruptly.
@@
We analyze the attention map at different points during training. We find that the attention map shows a sparse, interpretable pattern after the sudden loss drop, while no such pattern is shown before the sudden drop (Figure˜1). For the MWS task, this optimal attention pattern corresponds to each output token $y_{i}$ attending only to the input tokens relevant to its computation, i.e., attending to $x_{1}$ for $y_{1},$ and to $x_{i},x_{i-1}$ for $y_{i},i\geq 2.$ We further use an Attention Progress Measure (APM) to record the progress of the attention map toward its optimal pattern during training, defined as
${\rm APM}:=\frac{\sum_{(i,j)\in\Omega}|A_{ij}|}{\sum_{(i,j)}|A_{ij}|},$ (3)
where $A_{ij}$ denotes the attention score allocated to the $j^{th}$ token when computing output at the $i^{th}$ position in the sequence, and $\Omega$ is the set of position pairs in the optimal attention map. This measure is defined with absolute values due to our choice of linear attention so that $A_{ij}$ could be positive or negative. In experiments, we calculate $\rm APM$} averaged over a random batch of sequences. The APM sets $\Omega$ for all relevant tasks in this paper are defined in Table˜2.
Figure˜2(b) shows that the $\rm APM$} monotonically increases from near 0 to near $0.8$ during training, and its increase is more gradual than the loss/accuracy dynamics. In particular, $\rm APM$} already increases to a nontrivial value during the loss plateau and before the sudden loss drop.
@@ 3 Implicit Biases in the Early Phase of Training
During the loss plateau, the model often has already learned to implement a partial solution to the task. This means it correctly predicts a subset of the output tokens, typically those corresponding to an intuitively simpler part of the problem, while failing on the more complex parts. For instance, in the MWS task, the model quickly learns to predict the first output token $y_{1}$ correctly (see Figure˜2(a) for the first-token accuracy), as it is simply a copy of the first input token $x_{1}$, while the overall loss remains high and accuracy on subsequent tokens is poor. This ability to solve easier sub-components of the task early on is observed across various algorithmic problems (see Table˜1 in Appendix˜B).
@@
$\rho:=\frac{1}{n-1}\sum_{i=1}^{n-1}\mathbf{1}[y_{i}=y_{i+1}].$ (4)
We observe that $\rho$ increases rapidly during the early phase of training, when the optimal attention map has not been learned yet (Figure˜2(b)). Note that this frequency is small at initialization, and grows rapidly in the first $\approx 50$ steps to $\approx 0.8$, indicating that it is an implicit bias coming from gradient-based training.
@@
${\rm COS}_{i,j}:=\frac{\langle\mathbf{h}_{i},\mathbf{h}_{j}\rangle}{\|\mathbf{h}_{i}\|\|\mathbf{h}_{j}\|}$ (5)
where $\mathbf{h}_{i}\in\mathbb{R}^{d}$ is the hidden state at position $i$ in the sequence (this quantity is averaged over a random batch of sequences). We find that in the early phase of training, there is a rapid increase in ${\rm COS}_{i,j}$—averaged over all output positions $i,j$ except the first position, this quantity increases to $\approx 0.95$ (Figure˜2(b)). We emphasize that similar to repetitions, representation collapse is not present at initialization and only appears after a few steps of training. This is in contrast to the rank collapse phenomenon for deep softmax-attention Transformers [2] that occurs at initialization. Also, while we focus on the final-layer representation (before the LM layer) in the main text, we show in Figure˜42 that representation collapse happens in all intermediate layers to varying degrees.
@@ 4 The Role of Learning Attention
Observe that though the loss dynamics are abrupt, attention progress measure as well as repetitions and representation collapse are not (Figure˜2(b)); that is, even when the loss is barely decreasing (between steps 50 and 150), attention progress measure notably increases, accompanied by a decrease in repetition frequency and representation collapse. Via training-time interventions, this section shows that learning the attention map plays a crucial role in shaping the loss plateau as well as repetitions and representation collapse.
@@
We start by verifying whether the attention layer is responsible for representation collapse during the early phase of training. To this end, we plot the cosine similarity of the residual stream for output tokens just before and after the attention layer. Formally, let the residual stream before attention layer (i.e., token + positional embeddings) be $\mathbf{h}_{i}\in\mathbb{R}^{d}$, and the residual stream after attention layer be $\mathbf{h}^{\prime}_{i}\in\mathbb{R}^{d}$, we measure the norm and pairwise cosine similarity for $\mathbf{h}_{i}$ and $\mathbf{h}^{\prime}_{i}$ in Figure˜41.
@@
To study the role of attention map, we slightly modify the training process starting at different time points in training, biasing it towards (or away from) the optimal attention map to check if repetitions, representation collapse, and loss plateau are reduced (resp. amplified). We do the following: For each step in training after $t_{0}$, for attention map $A\in\mathbb{R}^{(L-1)\times(L-1)}$, we use the mask $M\in\mathbb{R}^{(L-1)\times(L-1)},M_{ij}=c$ for $(i,j)\in\Omega,M_{ij}=1$ otherwise, except for $i=1$ (since that is a partial solution and converges early in training). Then, we use the modified attention map $A\odot M$ (Hadamard product) for training and inference. Hence, for $c>1,$ this implies biasing the model towards the final (optimal) attention map, whereas for $0<c<1,$ this implies biasing the model away from the optimal attention map.
@@
We find that, for $c>1$ and various values of $t_{0},$ such a scaling leads to lower average cosine similarity between hidden states, lower frequency of repetitions, and faster convergence (Figures˜3 and 43). Whereas, for $0<c<1,$ we find the opposite: the model is in representation collapse state for a longer time and converges later compared to the non-scaled $(c=1)$ case, while the repetition frequency remains large throughout the plateau (Figure˜4).
For example, for $t_{0}=0,c=10$, i.e. scaling $10\times$ from the start of training, we find that the peak cosine similarity attained during training is $\approx 0.6$, much smaller than the $\approx 0.95$ attained for $c=1,$ and further the peak for $c=10$ is for negligible duration compared to that for $c=1$. Later values of $t_{0}=25,50,75$ show similar results wherein the cosine similarity drops immediately on the above biasing operation, followed by lower repetition frequency and convergence to optimal solution (Figure˜43). On the other hand, for $t_{0}=0,c=0.2,0.5$, the model takes much longer to converge and is in representation collapse / large repetition frequency state for much longer. This is in line with our expectation that lower attention map values for the optimal positions lead to slower learning and prolonged representation collapse.
@@
Training with Optimal Attention Map v. Other Components
In the first part of this test, we initialize with the optimal attention map by fixing embeddings, LayerNorm for attention layer and attention layer weights to their final values at the end of a normal training run, so that at initialization, the correct attention map is already available to subsequent layers. We re-train the subsequent non-fixed layers starting from random initialization. For the attention layer, we choose the set of parameters to initalize in 2 ways: (a) only Key, Query $(W_{K},W_{Q})$ weights, and (b) All of Key, Query, Value, Output $(W_{K},W_{Q},W_{V},W_{O})$ weights. We find that in both of these cases, learning only the subsequent layers (i.e. MLP, LM Head) take significantly shorter time than training the full model, without any significant representation collapse, repetitions or plateau in loss (Figure˜5). Further, between (a) and (b), we find that additionally having $W_{O},W_{V}$ layers initialized to optimal values slightly speeds up learning, and average cosine similarity goes up to approx $0.15$ instead of $\approx 0.45$ when only initializing $W_{K},W_{Q}$ weights. This indicates that $W_{O},W_{V}$ layers also play a non-trivial role in causing representation collapse.
On the other hand, fixing MLP or embeddings (together with LM head) to their final optimal values and re-training the other components does not qualitatively change the training dynamics from the full training case, i.e., a significant loss plateau, repetition bias, and representation collapse still occur (Figure˜5).
This result confirms that attention map is a major bottleneck that leads to early representation collapse and loss plateau, and that there is little benefit from having the optimal MLP or embeddings at initialization compared to attention map.
Our results are similar to [44] that shows how various sub-circuits (incl. attention map) for in-context learning affect the loss plateau via training-time interventions, and [29] that shows the effectiveness of attention transfer from pretrained model for downstream task training in Vision Transformers.
Tracking Progress in Hidden States
To further track model progress during training, we probe the output of Attention layer before it is added to the residual stream (Equation˜2) using a 2-layer MLP probe with ReLU activation, and find that the accuracy for this probe increases earlier than the model accuracy. Please see Appendix˜C for more details.
@@
Searching for the Right Tokens
The preceding results have demonstrated the important role of learning attention map (i.e., attending to the correct tokens) towards the plateau in training loss. Furthermore, once there is sufficient progress in the Attention map (visualized via APM in Figure˜2(b) and output probing in Figure˜20), the final model output converges rapidly to the correct solution. These observations conceptually align with abrupt learning observed when training 2–layer neural nets on a specific class of target functions [1], where the loss plateau is attributed to ‘search’ phase for aligning the first layer with the support of target function. In our setup, we hypothesize that a similar ‘search’ phenomenon could be occurring, where gradual increase in attention map weights at the required input tokens underlies the loss plateau.
Similar progress in model weights during loss plateau for a sinusoidal neuron trained on a sparse parity task was shown in [3, Fig. 3]. [38] discuss parallel search when training multi-head Transformers on a sparse parity task, and how the number of attention heads affects training dynamics.